Grade 9Mechanics

A body is moving with uniform acceleration covers 100m in the first 10s and 150m in the next 10s. Find the initial velocity of the body.

BHAVIN
9 Years agoGrade 9

2 Answers

Vikas TU
9 Years ago
From the second law of Newton’s eqn.,
100 = 10u+50a;
50 = 10u+ 40a;
50 = 10a;
a = 5 m/sec2;
150 = 10 u;
u = 15 m/sec
govind
7 Years ago
Let initial velocity =u m/s, acceleration = a m/s² final velocity ( the velocity at the end of first 10 seconds) = v m/s
Now for fist case time t = 10 sec.
Using second equation of motion s = ut + (1/2) at² we get
100 = u x 10 + (1/2) a (10)² = 100 = u x 10 + (1/2 ) x a x 10 x 10
= 100 = 10{u + (1/2) x a x 10) = 100/10 = u + 5a ( rearranging equation)
= u + 5a = 10 .........(1)
Now for second case there is one important point to consider that the initial velocity for the second case will be the final velocity of the first case as this motion starts after the end of first motion.
Thus to find initial velocity for second case we use first equation of motion v = u + at on first motion we get
v = u + a x 10 = v = u + 10a
Now this will be our initial velocity for second motion
Using second equation of motion s = ut + (1/2) gt² for second motion we get
150 = (u + 10a)x10 + (1/2) a (10)² = 150 = 10x(u + 10a) + (1/2) g x10x10
= 150 = 10 {(u +10a) + (1/2) x 10 x a} = 150/10 = u + 10a + 5a (rearranging equation)
= 15 = u + 15a= u + 15a = 15 ......(2)
Now subtracting equation (1) from equation (2) we get
u + 15a - (u + 5a) = 15 - 10 = u + 15a - u - 5a = 5
= u - u + 15a - 5a = 5 (rearranging equation)
= 10a = 5 = a = 5/10
= a = 1/2
Putting this value of a in equation (1) we get
u + 5 x (1/2) = 10 = u + (5/2) = 10
= u = 10 - 5/2 = 10 - 2.5
= u = 7.5 m/s